# Lossless data compression at pragmatic rates

### v3

@@ 3.2 Converse
Consider a memoryless source $X$ with p.m.f. $P$ with full support on $A$. Let $\delta\in(0,D(U\|P))$, where $U$ is the uniform distribution over $A$. Then for any $n>N_{0}$,
$R_{n}^{*}(2^{-n\delta})\geq H(P_{\alpha^{*}})-\frac{1}{2(1-\alpha^{*})}\frac{\log n}{n}-\frac{C}{n},$ (16)
@@
Proof. Taking $k=nR$, for some $R>0$, letting $\tau=n\delta$, and replacing $X$ with $X_{1}^{n}$ (and $f^{*}$ with $f_{n}^{*}$) in the one-shot converse in Theorem 2.3, yields:
$\mathbb{P}\Big{[}\ell(f_{n}^{*}(X_{1}^{n}))\geq n\Big{(}R-\frac{1}{n}\Big{)}\Big{]}>\mathbb{P}[-\log P^{n}(X_{1}^{n})\geq n(R+\delta)]-2^{-n\delta}.$
Therefore, if we can find a rate $R$ for which,
$\mathbb{P}[\log P^{n}(X_{1}^{n})\leq-n(R+\delta)]\geq 2\times 2^{-n\delta},$
this would imply that $\mathbb{P}[\ell(f_{n}^{*}(X_{1}^{n}))\geq n(R-1/n)]>2^{-n\delta}$, which would in turn give,
$R_{n}^{*}(2^{-n\delta})\geq R-\frac{1}{n}.$ (17)
Writing,
$S_{n}=\{x_{1}^{n}\in A^{n}:\log P^{n}(x_{1}^{n})\leq-n(R+\delta)\},$
our objective then is to find an appropriate rate $R$ for which,
@@
First we will use Lemma 2.4 to express $P^{n}(S_{n})$ in terms of $P_{\alpha_{n}}^{n}(S_{n}^{c})$ for some $\alpha_{n}\in(0,1)$ that will be chosen later. Indeed, since,
$S_{n}$ =$=\Bigg{\{}x_{1}^{n}\in A^{n}:\sum_{i=1}^{n}\log P^{1-\alpha_{n}}(x_{i})\leq-(1-\alpha_{n})n(R+\delta)\Bigg{\}}$
=$=\Bigg{\{}x_{1}^{n}\in A^{n}:\log\frac{P^{n}(x_{1}^{n})}{P^{n}_{\alpha_{n}}(x_{1}^{n})}\leq-(1-\alpha_{n})n(R+\delta)+n\log Z_{\alpha_{n}}\Bigg{\}},$ (19)
we have that,
$\frac{P^{n}(x_{1}^{n})}{P_{\alpha_{n}}^{n}(x_{1}^{n})}\geq\max_{x_{1}^{n}\in S_{n}}\frac{P^{n}(x_{1}^{n})}{P_{\alpha_{n}}^{n}(x_{1}^{n})}\quad\mbox{for all}\;x_{1}^{n}\not\in S_{n},$
so we can apply Lemma 2.4 to obtain that,
@@
The next step will be an application of the Berry-Esséen bound to obtain an upper bound on $P_{\alpha_{n}}^{n}(S_{n}^{c})$. From (19) we get,
@@
and the Berry-Esséen in Theorem 2.1 applied to (21) yields,
@@
Using Lemma 2.6, the bound becomes,
@@
Using this bound, from (20) we get,
@@
Then, using the hypothesis testing converse in Theorem 2.5, with $P_{\alpha_{n}}$ in the place of $Q$, $\epsilon=\frac{1}{2}+\frac{\hat{\rho_{3}}}{2\tilde{\sigma_{3}}^{3}\sqrt{n}}$, and $\Delta=1$, yields,
$\log P^{n}(S_{n})$ $\geq-nD(P_{\alpha_{n}}\|P)-\sigma_{2,\alpha_{n}}(\log e)\sqrt{n}\Phi^{-1}\Big{(}\frac{1}{2}+\frac{\hat{\rho_{3}}}{2\tilde{\sigma_{3}}^{3}\sqrt{n}}+\frac{\rho_{2,\alpha_{n}}}{2\sigma_{2,\alpha_{n}}^{3}\sqrt{n}}+\frac{1}{\sqrt{n}}\Big{)}-\frac{1}{2}\log n,$ (24)
@@
Recall that the existence of a unique $\alpha^{*}$ as in the statement of the theorem is already established in Theorem 3.1. In order to bound $\log P^{n}(S_{n})$ below further, we will use simple Taylor expansions to get bounds on $D(P_{\alpha_{n}}\|P)$, $\sigma_{2,\alpha_{n}}$ and the $\Phi^{-1}$ term in (25). First we expand $D(P_{\alpha_{n}}\|P)$ around $\alpha^{*}$. Using the expressions for the derivatives of $D(P_{\alpha}\|P)$ in Lemma 2.7, gives,
@@
for some $u\in[\alpha^{*},\alpha_{n}]$, assuming $\alpha^{*}\leq\alpha_{n}$, and where $\sigma_{3}=\sigma_{3,\alpha^{*}}$. Letting $\hat{\sigma}_{3}^{2}=\sup_{\alpha\in(0,1)}\sigma_{3,\alpha}^{2}$, we have $\sigma_{3,u}^{2}\leq\hat{\sigma}_{3}^{2}$, and we observe that $|u-1|\leq 1$ and that,
@@
Next, since $\sigma_{2,\alpha_{n}}=(1-\alpha_{n})\sigma_{3,\alpha_{n}}$ by Lemma 2.6, we can expand $\sigma_{3,\alpha_{n}}$ instead of $\sigma_{2,\alpha_{n}}$ around $\alpha^{*}$, yielding,
@@
And for the third term, we expand $\Phi^{-1}$ around $1/2$ to obtain that,
@@
Finally, substituting the bounds (27), (28) and (29) into (25), gives,
@@
Now we are in a position to choose $\alpha_{n}$. Letting,
@@
simple algebra shows that we have,
@@
as long as $n$ is large enough so that,
@@
To that end, starting from the definition of $R$ in (22), we can re-write,
@@
To bound this below, we will expand $D(P_{\alpha_{n}}\|P)$ and $H(P_{\alpha_{n}})$ around $\alpha^{*}$. For $D(P_{\alpha_{n}}\|P)$, continuing from the Taylor expansion in (26) we have,
@@
where (37) holds for some $v\in[\alpha^{*},\alpha_{n}]$; (38) holds since $\sigma_{3,u}^{2}\leq\hat{\sigma_{3}}^{2}$, $v\leq 1$ and by the definition of $\hat{\rho_{3}}$; (39) follows from the definition of $\alpha_{n}$ and the fact that $n$ is taken sufficiently large as above; and (40) follows by substituting the expression of $\alpha_{n}$ in (31).
Substituting the bounds (36) and (40) into (33), gives,

### v4

@@ 3.2 Converse
[…]
$R_{n}^{*}(2^{-n\delta},P)\geq H(P_{\alpha^{*}})-\frac{1}{2(1-\alpha^{*})}\frac{\log n}{n}-\frac{C}{n},$ (18)
@@
Proof. Taking $k=nR$, for some $R>0$, letting $\tau=n\delta$, and replacing $X$ with $X^{n}$ (and $f^{*}$ with $f_{n}^{*}$) in the one-shot converse in Theorem 2.3, yields:
$\mathbb{P}\Big[\ell(f_{n}^{*}(X^{n}))\geq n\Big(R-\frac{1}{n}\Big)\Big]>\mathbb{P}[-\log P^{n}(X^{n})\geq n(R+\delta)]-2^{-n\delta}.$
Therefore, if we can find a rate $R$ for which
$\mathbb{P}[\log P^{n}(X^{n})\leq-n(R+\delta)]\geq 2\times 2^{-n\delta},$
this would imply that $\mathbb{P}[\ell(f_{n}^{*}(X^{n}))\geq n(R-1/n)]>2^{-n\delta}$, which would in turn give
$R_{n}^{*}(2^{-n\delta},P)\geq R-\frac{1}{n}.$ (19)
Writing
$S_{n}=\{x^{n}\in A^{n}:\log P^{n}(x^{n})\leq-n(R+\delta)\},$
our objective then is to find an appropriate rate $R$ for which
@@
First we will use Lemma 2.4 to express $P^{n}(S_{n})$ in terms of $P_{\alpha_{n}}^{n}(S_{n}^{c})$ for some $\alpha_{n}\in(0,1)$ that will be chosen later. Indeed, since
$S_{n}$ $=\Bigg\{x^{n}\in A^{n}:\sum_{i=1}^{n}\log P^{1-\alpha_{n}}(x_{i})\leq-(1-\alpha_{n})n(R+\delta)\Bigg\}$
$=\Bigg\{x^{n}\in A^{n}:\log\frac{P^{n}(x^{n})}{P^{n}_{\alpha_{n}}(x^{n})}\leq-(1-\alpha_{n})n(R+\delta)+n\log Z_{\alpha_{n}}\Bigg\},$ (21)
we have that
$\frac{P^{n}(x^{n})}{P_{\alpha_{n}}^{n}(x^{n})}\geq\max_{x^{n}\in S_{n}}\frac{P^{n}(x^{n})}{P_{\alpha_{n}}^{n}(x^{n})}\quad\mbox{for all}\;x^{n}\not\in S_{n},$
so we can apply Lemma 2.4 to obtain
@@
The next step will be an application of the Berry-Esséen bound to obtain an upper bound on $P_{\alpha_{n}}^{n}(S_{n}^{c})$. From (21) we get
@@
and the Berry-Esséen in Theorem 2.1 applied to (23) yields
@@
Using Lemma 2.6, the bound becomes
@@
Using this bound, from (22) we get
@@
Then, using the hypothesis testing converse in Theorem 2.5, with $P_{\alpha_{n}}$ in the place of $Q$, $\epsilon=\frac{1}{2}+\frac{\hat{\rho}_{3}}{2\tilde{\sigma}_{3}^{3}\sqrt{n}}$, and $\Delta=1$, yields
$\log P^{n}(S_{n})$ $\geq-nD(P_{\alpha_{n}}\|P)-\sigma_{2,\alpha_{n}}(\log e)\sqrt{n}\Phi^{-1}\Big(\frac{1}{2}+\frac{\hat{\rho}_{3}}{2\tilde{\sigma}_{3}^{3}\sqrt{n}}+\frac{\rho_{2,\alpha_{n}}}{2\sigma_{2,\alpha_{n}}^{3}\sqrt{n}}+\frac{1}{\sqrt{n}}\Big)-\frac{1}{2}\log n$ (26)
@@
Recall that the existence of a unique $\alpha^{*}$ as in the statement of the theorem is already established in Theorem 3.1. In order to bound $\log P^{n}(S_{n})$ below further, we will use simple Taylor expansions to get bounds on $D(P_{\alpha_{n}}\|P)$, $\sigma_{2,\alpha_{n}}$ and the $\Phi^{-1}$ term in (27). First we expand $D(P_{\alpha_{n}}\|P)$ around $\alpha^{*}$. Using the expressions for the derivatives of $D(P_{\alpha}\|P)$ in Lemma 2.7, gives
@@
for some $u\in[\alpha^{*},\alpha_{n}]$, assuming $\alpha^{*}\leq\alpha_{n}$, and where $\sigma_{3}=\sigma_{3,\alpha^{*}}$. Letting $\hat{\sigma}_{3}^{2}=\sup_{\alpha\in(0,1)}\sigma_{3,\alpha}^{2}$, we have $\sigma_{3,u}^{2}\leq\hat{\sigma}_{3}^{2}$, and we observe that $|u-1|\leq 1$ and that
@@
Next, since $\sigma_{2,\alpha_{n}}=(1-\alpha_{n})\sigma_{3,\alpha_{n}}$ by Lemma 2.6, we can expand $\sigma_{3,\alpha_{n}}$ instead of $\sigma_{2,\alpha_{n}}$ around $\alpha^{*}$, yielding
@@
And for the third term, we expand $\Phi^{-1}$ around $1/2$ to obtain that
@@
Finally, substituting the bounds (29), (30) and (31) into (27), gives
@@
Now we are in a position to choose $\alpha_{n}$. Letting
@@
simple algebra shows that we have
@@
as long as $n$ is large enough so that
@@
To that end, starting from the definition of $R$ in (24), we can re-write
@@
To bound this below, we will expand $D(P_{\alpha_{n}}\|P)$ and $H(P_{\alpha_{n}})$ around $\alpha^{*}$. For $D(P_{\alpha_{n}}\|P)$, continuing from the Taylor expansion in (28) we have
@@
where (39) holds for some $v\in[\alpha^{*},\alpha_{n}]$; (40) holds since $\sigma_{3,v}^{2}\leq\hat{\sigma}_{3}^{2}$, $v\leq 1$ and by the definition of $\hat{\rho}_{3}$; (41) follows from the definition of $\alpha_{n}$ and the fact that $n$ is taken sufficiently large as above; and (42) follows by substituting the expression of $\alpha_{n}$ in (33).
Substituting the bounds (38) and (42) into (35), gives